# EquiBot: SIM(3)-Equivariant Diffusion Policy for Generalizable and Data Efficient Learning

### v1

@@ 4 Experiments
Through our experiments, we want to answer the following questions: (1) does our method generalize to unseen scenarios better than imitation learning methods that do not leverage equivariance; (2) does our method demonstrate more robust performance than prior methods for equivariant visuomotor policy learning that do not leverage diffusion models; (3) does our method achieve better data efficiency when there is little training data compared to prior imitation learning methods? To answer these three questions, we perform quantitative experiments in both simulation (Section 4.1) and the real world (Section 4.2).
@@ 4.1.1 Comparisons to Vanilla Diffusion Policies and Other Equivariant Policy Architectures
Comparisons. We compare our method to three baselines. (1) Diffusion Policy (DP) [1]: vanilla diffusion policy modified to use point cloud inputs; we substitute the image-based encoder with a PointNet++ encoder [32]. (2) Diffusion Policy with Augmentations (DP+Aug): this baseline uses the same architecture as the vanilla diffusion policy baseline, but trains with synthetically generated data augmentation. (3) EquivAct [21]: a re-implementation of [21] that drops the pre-training phase that requires task-specific simulated data.
@@
Tasks. We use four simulated tasks: Cloth Folding, Object Covering, Box Closing, and Push T (see Figure 3). The first three tasks involve two mobile robots manipulating various deformable and articulated objects. In these tasks, a simulated depth camera records point clouds of relevant objects in the scene from a third-person viewpoint. The policy takes these point clouds as input and commands the end-effector position, rotation, and gripper open-close actions of both robots. The Push T benchmark task is a simulated 2D T-shape pushing game developed in [1] to showcase learning from multimodal demonstrations. To make this task compatible with our setup, we assume the robot and object are placed on the ground plane ($z=0$) in 3D space. In this task, the policy receives the positions of the eight corners of the T-shape as input and outputs a position target to the pushing operator.
@@
Augmentations. The DP+Aug baseline requires augmentations in the training phase. In all four environments, we augment the training data to (1) rotate the observation up to 360 degrees around the z-axis, (2) uniformly scale the observation within the range 0.5×0.5$\times$0.5 × to 1.5×1.5$\times$1.5 ×, and (3) apply a random Gaussian offset to the observation with standard deviation equal to $0.1$ times the approximate workspace size.
Training and evaluation. We train all methods for 2,000 epochs. For every training run, we save a checkpoint every 50 epochs and evaluate the last 5 checkpoints saved at the end of training. For every evaluation, we run the policy in a randomly initialized environment for 10 episodes and record the mean final reward achieved by the policy. For all results we show, we run training over 3 random seeds and report the mean and standard deviation of evaluation results over these seeds.
@@
Figure 4: Results of out-of-distribution generalization experiments. We show that our method achieves more robust out-of-distribution generalization performance than methods that do not use diffusion processes to model policies and ones that do not utilize equivariance. Error bars show the mean and standard deviation over 3 seeds.
@@
Evaluation setups. To gain insight into the generalizability of competing methods, we design four different evaluation setups to test our trained policies. The Original setup evaluates the trained policy at the exact same initial poses and goals as in the demos; the OOD (R+Su) setup randomizes initial object rotation with randomized object uniform scaling from 1×1$\times$1 × to 2×2$\times$2 ×; the OOD (R+Sn) setup randomizes rotation and scaling as in OOD (R+Su), but additionally adds non-uniform scaling up to a $1.33$ aspect ratio change; the OOD (R+Sn+P) setup adds large position randomization to the OOD (R+Sn) setup.
@@
Results. We show the results of this experiment in Figure 4. The DP baseline performs very well in the Original setup, but its performance drops significantly when it comes to any of the OOD setups. The DP+Aug baseline has slightly worse performance compared to DP in the Original setup. This is because DP+Aug trains not only from original data but also from more synthetically augmented scenarios. Thus, DP+Aug likely allocates less representational capacity to the original in-distribution data. When it comes to OOD setups, DP+Aug performs much better than DP, but still suffers from performance drops. The EquivAct baseline performs very well in the Cloth Folding task but displays subpar performance in Object Covering and Box Closing tasks due to unstable training performance among checkpoints. It also does not perform well in the Push T task because it cannot handle multi-modal training data well. Our method performs stably in all four tasks and suffers from the least amount of performance drop compared to all baselines. This shows that our method indeed outperforms prior methods in out-of-distribution generalization.
@@ 4.1.2 Data Efficiency Experiments
Figure 5: Results of data efficiency experiments. Our method achieves better data efficiency than diffusion policy when evaluated in-distribution on two benchmark tasks.
In this experiment, we aim to test if our method outperforms prior methods in a low-data regime, even during in-distribution evaluation. Because we only care about in-distribution performance in this experiment, we only compare our method with the DP baseline. We adopt two Robomimic environments [34] for this experiment: Can and Square.
Setup. In each task, we train all methods in three setups: learning from 100 demos, 50 demos, and 25 demos. We run 2,000 epochs for each method with 3 random seeds. As in the previous experiment, each evaluation job computes the average performance over the last five checkpoints in the training run with 50-epoch intervals. The results plot the mean and standard deviation over seeded runs.
Results. The results of this experiment can be found in Figure 5. In both tasks, the performance of DP drops dramatically when the number of demos decreases from 100 to 25, while our method retains relatively higher performance when the amount of data drops. This is because when training data size is small, the training data does not cover the whole distribution of initial poses to allow the DP baseline to learn how to complete the task with all possible initial poses during evaluation. Our method, on the other hand, leverages equivariance to generalize to varied initial object poses better than the DP baseline.
@@ 4.2 Real Robot Experiments
Figure 6: Real robot evaluation setups. Each block represents one task. In each block, we show sample training demonstrations collected by a human on the left and evaluation scenarios on the right. We show qualitative results for more tasks in the supplementary materials (Section B.2).
Push Chair Luggage Packing Luggage Closing Unseen Poses $\rightarrow$ Translate + Rotate Object Variations $\rightarrow$ Long Desk Round Table T-shirt Towel Roll Cap Shorts Large Luggage DP 0/10 0/10 0/10 0/10 0/10 0/10 0/10 DP+Aug 0/10 0/10 0/10 0/10 0/10 0/10 2/10 Ours 8/10 10/10 7/10 3/10 8/10 8/10 6/10
Laundry Door Closing Bimanual Folding Bimanual Make Bed Unseen Poses $\rightarrow$ Translate + Rotate Object Variations $\rightarrow$ Long Bath Towel Comforter DP 3/10 0/10 0/10 DP+Aug 1/10 0/10 0/10 Ours 8/10 6/10 8/10
@@
In our real robot experiments, we show a series of experiments where we train mobile robots to perform everyday manipulation tasks from 5 minutes of single-view human demonstration videos. We select a suite of 6 tasks that involve diverse everyday objects, including rigid, articulated, and deformable objects (see Figure 6): (1) Push Chair: A robot pushes a chair towards a desk; (2) Luggage Packing: A robot picks up a pack of clothes and places it inside an open suitcase; (3) Luggage Closing: A robot closes an open suitcase on the floor; (4) Laundry Door Closing: A robot pushes the door of a laundry machine to close it; (5) Bimanual Folding: Two robots collaboratively fold a piece of cloth on a couch; (6) Bimanual Make Bed: Two robots unfold a comforter to make it cover the bed completely.
Data collection. We collect 15 human demonstration videos for each real robot task. We use a ZED 2 stereo camera to record the movement of a human operator using their fingers to manipulate the objects of interest at 15 Hz. After data collection, we use an off-the-shelf hand detection model, an object segmentation model, and a stereo-to-depth model to parse out the human hand poses and object point clouds in each frame of the collected demos. We then subsample this data to 3 Hz and convert it into a format supported by our policy training algorithm. More details regarding our data collection and demonstration processing system can be found in the supplementary materials (Section E.1).
Mobile robot setup. In all real robot experiments, we use holonomic mobile bases with Kinova Gen3 7 DoF arms mounted on top. We use a single ZED 2 camera mounted at a fixed position in the workspace to obtain visual observations for the robot policy. We first use the ZED 2 camera to obtain a single-view point cloud of the scene, and then use the Grounded Segment Anything Model [35] to segment out the relevant objects in the scene based on a natural language description of objects involved in the task. The segmented single view point cloud is then used as the input visual observation for the policy.
We utilize the motion capture system in the room to track the pose of the mobile base. During one evaluation, the learned policy reads parsed point clouds and sends 8 actions at a time to the robot to execute. The mobile robot then moves the arm to achieve the commanded actions, moving the mobile base when the arm is too close or too far away from the base. Please see supplementary materials (Section E.2) for more details.
@@
Training and evaluation. We train all methods for 1,000 epochs, then evaluate each method for 10 episodes and record the success rate. We vary the evaluation scenarios from the training scenarios differently in each task. In Laundry Door Closing, we perform evaluations in-distribution. In Push Chair, Luggage Closing, and Bimanual Make Bed, we evaluate with novel objects to make the evaluation out-of-distribution to the training data. In Luggage Packing and Bimanual Folding, we not only switch to novel objects but also translate and rotate the layout of the scene.
@@
Results. The results of real robot experiments are shown in Table 1. The evaluation shows that our method can generalize to diverse unseen objects, outperforming the DP baseline in both in-distribution and out-of-distribution scenarios with novel objects and unseen object poses. See supplementary materials for more results (Section B.2) and detailed analysis (Section B).
5 Conclusion
@@
We proposed EquiBot, a visuomotor policy learning method that is capable of generalizable and data-efficient policy learning in a wide range of robot manipulation tasks. In a suite of 6 simulated and 6 real robot tasks, we showed that our proposed method outperforms vanilla diffusion policies and prior imitation learning methods using equivariant architectures. We demonstrated that our method can learn from just 5 minutes of human demonstrations and generalize to unseen scenarios that are dramatically different from training scenarios.
5.1 Limitations and future work
While our method generalizes to scenes with unseen object positions, scales, and orientations, it does not handle nonlinear changes in object shapes or dynamics by construction. For example, a change in object shape beyond re-scaling: a kettle with a very short spout in demonstrations versus a very long curved spout at deployment. An example with a change in dynamics: demonstrations with a stiff cloth versus using cloths that stretch or sag significantly more at deployment. Extending our method to multi-task setups and training to handle such nonlinear data transformations would be an interesting direction for future work.
@@ Acknowledgments
This work was supported in part by the Toyota Research Institute. Stanford Robotics Center provided space for our experiments.
@@
Supplementary Materials
Supplementary Materials A Further Discussion of Related Work
Imitation learning with diffusion policies from 3D inputs. While many imitation learning works assume RGB images as visual observations [34, 2, 7, 1], some works, similar to the design choice of our work, assume 3D point clouds or depth inputs to their methods. Recently, Ze et al. [29] proposed 3D Diffusion Policy, a depth-only variant of diffusion policy for visuomotor policy learning. Their method is very similar to our DP baseline, with two differences: (1) they use a simpler DP3 encoder in their work; (2) they use a 2-layer MLP to encode the robot proprioceptive states before concatenating with the point cloud representation.
@@
We also show that [29] can be easily integrated with our method by switching our PointNet-based encoder to the DP3 encoder. In Figure 7, we show a comparison between our method and a variation of our method with a modification of the DP3 encoder to make it SIM(3)-equivariant. Results show that the DP3 variant has slightly lower but comparable performance as our method in the Cloth Folding task.
Supplementary Materials B Further Evaluations and Performance Analysis
B.1 Failure Analysis

### v2

@@ 4 Experiments
Through our experiments, we want to answer the following questions: (1) does our method generalize to unseen scenarios better than imitation learning methods that do not leverage equivariance; (2) does our method demonstrate more robust performance than prior methods for equivariant visuomotor policy learning that do not leverage diffusion models; (3) does our method achieve better data efficiency when there is limited training data; (4) how do different components of the method contribute to the final performance? We perform quantitative experiments in both simulations (Section 4.1) and the real world (Section 4.2) to answer these questions. Please view ablation experiments in Section B of the supplementary materials.
@@ 4.1.1 Comparisons to Vanilla Diffusion Policies and Other Equivariant Policy Architectures
Comparisons. We compare our method to three baselines. (1) Diffusion Policy (DP) [8]: Vanilla diffusion policy using a point cloud as input; we substitute the imaged-based encoder to a PointNet++ encoder [39] similar to what EquiBot is using. (2) Diffusion Policy with Augmentations (DP+Aug): This baseline uses the same architecture as the vanilla diffusion policy baseline, but trains with synthetically generated data augmentation. (3) EquivAct [16]: A re-implementation of [16] that drops the pre-training phase that requires task-specific simulated data.
@@
Augmentations. The DP+Aug baseline requires augmentations in the training phase. In all four environments, we augment the training data to (1) rotate the observation around the z-axis, (2) uniformly scale the observation within the range 0.5×1.5×0.5$\times$-1.5$\times$0.5 × - 1.5 ×, and (3) apply a random Gaussian offset to the observation with standard deviation equal to $0.1$ times the approximate workspace size.
@@
Figure 4: Results of out-of-distribution generalization experiments. We show that our method achieves more robust out-of-distribution generalization performance than methods that do not use diffusion processes to model policies and ones that do not utilize equivariance. Error bars show the mean and standard deviation over 5 checkpoints and 3 seeds.
Tasks. We use four simulated tasks: Cloth Folding, Object Covering, Box Closing, and Push T (see Figure 3). The first three tasks involve two mobile robots manipulating various deformable and articulated objects. In these tasks, a simulated depth camera records point clouds of relevant objects in the scene from a third-person viewpoint. The policy takes these point clouds as input and commands the end-effector position, rotation, and gripper open-close actions of both robots. We train policies in these tasks with 50 synthetically generated demonstrations. The Push T benchmark task is a simulated 2D T-shape pushing game developed in [8] to showcase learning from multimodal demonstrations. To make this task compatible with our setup, we assume the agent and object are placed on the ground plane ($z=0$) in 3D space. In this task, the policy receives the eight corners of the T-shape as input and outputs the velocity command to the pushing operator. We use the same 200 demonstrations as [8] to train policies in this task.
Training and evaluation. We train all methods for 2,000 epochs on 3 random seeds. For every training run, we save a checkpoint every 50 epochs and evaluate the last 5 checkpoints saved at the end of training. For each evaluation, we run the policy in a randomly initialized environment for 10 episodes and record the mean final reward the policy achieves. This means that each bar in the resulting plot reports the mean and standard deviation of evaluation results over $3$ seeds ×$\times 5$ last checkpoints ×$\times 10$ episodes $=150$ trials of evaluation.
@@
Evaluation setups. To gain insight into the generalizability of competing methods, we design four different evaluation setups to test our policies. The Original setup evaluates the policy at the same initial poses and goals as in the demos; the OOD (R) setup randomizes initial object rotation around the z-axis; the OOD (S) setup scales the scene up 1×1$\times$1 × to 2×2$\times$2 × with up to a $1.33$ aspect ratio change; the OOD (P) setup adds dramatic position randomization to the scene; the OOD (R+S+P) combines previous OOD setups by randomizing rotation, scale, and translation of the scene.
@@
Results. We show the results of this experiment in Figure 4. The DP baseline performs very well in the Original setup, but its performance drops significantly when it comes to any of the OOD setups. The DP+Aug baseline performs especially well in the OOD (R) setup because the augmentations data this baseline learns from are in-distribution with respect to the evaluation scenarios. But in other scenarios such as OOD (P) and OOD (R+S+P), it suffers from significant performance drops. The EquivAct baseline performs very well in the Cloth Folding task but displays subpar performance in Object Covering and Box Closing tasks. It also cannot perform well in the Push T task because the deterministic network architecture with behavior cloning loss cannot handle multi-modal training data well. We also show in supplementary materials Section B that EquivAct displays unstable training performance. Our method performs stably in all four tasks and suffers from the least amount of performance drop compared to all baselines. This shows that our method indeed outperforms prior methods in out-of-distribution generalization.
@@ 4.1.2 Data Efficiency Experiments
Figure 5: Results of data efficiency experiments. Our method achieves better data efficiency than the Diffusion Policy when evaluated in distribution on two benchmark tasks.
In this experiment, we aim to test if our method outperforms prior methods in a low-data regime, even if it is evaluated in distribution. Because we only care about in-distribution performance in this experiment, we only compare our method with the DP baseline. We adopt two Robomimic environments [15] for this experiment: Can and Square.
Setup. In each task, we train all methods 2,000 epochs in three setups: learning from 100 demos, 50 demos, and 25 demos. Following the same evaluation standard for mobile dual-robot and Push T environments, we report the mean and standard deviation over 150 trials of evaluation.
Results. The results of this experiment can be found in Figure 5. In both tasks, the performance of DP drops dramatically when the number of demos decreases from 100 to 25, while our method retains relatively higher performance when the amount of data drops. This is because when the training data size is too small to cover the whole distribution of initial poses, the DP baseline cannot naturally generalize to unseen initial poses during evaluation. Our method, on the other hand, leverages the equivariance nature to cope with novel initial object poses at test time.
@@ 4.2 Real Robot Experiments
We show a series of real robot experiments where we train mobile robots to perform everyday manipulation tasks from 5 minutes of single-view human demonstration videos. We select a suite of 6 tasks that involve diverse everyday objects, including rigid, articulated, and deformable objects (see Figure 6): (1) Push Chair: A robot pushes a chair towards a desk; (2) Luggage Packing: A robot picks up a pack of clothes and places it in an open suitcase; (3) Luggage Closing: A robot closes an open suitcase on the floor; (4) Laundry Door Closing: A robot pushes the door of a laundry machine to close it; (5) Bimanual Folding: Two robots collaboratively fold a piece of cloth on a couch; (6) Bimanual Make Bed: Two robots unfold a comforter to make it cover the bed completely.
Data collection and robot setup. We collect 15 human demonstration videos for each real robot task. We use a ZED 2 stereo camera to record the movement of a human operator using their fingers to manipulate the objects of interest at 15 Hz. After data collection, we use an off-the-shelf hand detection model [41], an object segmentation model [42], and a proprietary learned stereo-to-depth model to parse out the human hand poses and object point clouds in each frame of the collected demos. We then subsample this data to 3 Hz and convert it into a format supported by our policy training algorithm. In all real robot experiments, we use holonomic mobile bases [43] with Kinova Gen3 7 DoF arms mounted on top. Similar to human demonstration processing, we use a ZED2 camera and an object segmentation model to obtain the processed segmented point cloud as part of the input of our policy. See Section F of supplementary materials for more details.
Figure 6: Real robot evaluation setups. Each block represents one task. In each block, we show sample training demos collected by a human on the left and evaluation scenarios on the right.
Push Chair Laundry Door Closing Luggage Closing Unseen Poses $\rightarrow$ Object Variations $\rightarrow$ Long Desk Round Table Large Luggage DP 0/10 0/10 3/10 0/10 DP+Aug 1/10 2/10 2/10 2/10 Ours 8/10 10/10 8/10 6/10
Luggage Closing Bimanual Folding Bimanual Make Bed Unseen Poses $\rightarrow$ Translate + Rotate Translate + Rotate Object Variations $\rightarrow$ T-shirt Towel Roll Cap Shorts Long Bath Towel Comforter DP 0/10 0/10 0/10 0/10 0/10 0/10 Ours 7/10 3/10 8/10 8/10 6/10 8/10
@@
Training and evaluation. We train all methods for 1,000 epochs. After training, we evaluate each method for 10 episodes and record the success rate of the method. We vary the evaluation scenarios from the training scenarios differently in each task. In Laundry Door Closing, we perform evaluations in-distribution. In Push Chair, Luggage Closing, and Bimanual Make Bed, we evaluate with out-of-distribution objects. In Luggage Packing and Bimanual Folding, we not only switch to novel objects but also translate and rotate the layout of the scene.
@@
Results. The results of real robot experiments are shown in Table 1. The evaluation shows that our method can generalize to diverse unseen objects, outperforming the DP baseline in both in-distribution and out-of-distribution scenarios with novel objects and unseen object poses. See supplementary materials Section C for more results and detailed analysis.
5 Conclusions, Limitations and Future work
@@
We proposed EquiBot, a visuomotor policy learning method for generalizable and data-efficient policy learning in a wide range of robot manipulation tasks. In a suite of 6 simulated and 6 real robot tasks, we showed that our proposed method outperforms vanilla diffusion policies and prior imitation learning methods using equivariant architectures. We demonstrated that our method can learn from just 5 minutes of human demonstrations and generalize to unseen scenarios that are dramatically different from training scenarios.
While our method generalizes to scenes with unseen object positions, scales, and orientations, it does not handle nonlinear changes in object shapes or dynamics by construction. Our method also does not handle variations in the relative positioning of objects when multiple objects are present. Resolving these limitations might involve explicitly modeling scene dynamics and individual objects. Our method might also fail when the scene is partially occluded or the camera angle changes dramatically. This issue can be solved by learning general-purpose 3D representations robust to incomplete point clouds or novel viewing angles. Please see more discussions on limitations in Section H of supplementary materials. Solving the above challenges in out-of-distribution generalization and extending to multi-task setups are interesting directions for future work.
@@ Acknowledgments
This work was supported in part by the Toyota Research Institute, Intrinsic and the Human-Centered AI Institute. The Stanford Robotics Center provided the space for our experiments.
@@
Appendix A Further Discussion of Related Works
Imitation learning from 3D inputs. While many imitation learning works assume RGB images as visual observations [15, 4, 20, 8], some works, similar to the design choice of our work, assume 3D point clouds or depth inputs to their methods. Recently, Ze et al. [44] proposed 3D Diffusion Policy, a depth-only variant of diffusion policy for visuomotor policy learning. Their method is very similar to our DP baseline, with two differences: (1) they use a simpler DP3 encoder in their work; (2) they use a 2-layer MLP to encode the robot proprioceptive states before concatenating with the point cloud representation.
@@ Appendix A Further Discussion of Related Works
We also show that [44] can be easily integrated with our method by switching our PointNet-based encoder to the DP3 encoder. In Figure 7, we show a comparison between our method and a variation of our method with a modification of the DP3 encoder to make it SO(3)-equivariant. Results show that the DP3 variant has slightly lower but comparable performance as our method in the Cloth Folding task.
Figure 8: Ablations on model architecture. EquiBot outperforms naive extensions of EquivAct [16] for a variety of hyperparameter choices.
Additional works in equivariant architectures for robot manipulation. Aside from prior equivariant architectures for robot manipulation introduced in the main paper, there are also robotics works that attempt to utilize equivariance in various setups. Some works [45, 46] attempt to use equivariance in a pick-and-place setup, while others [47, 48] propose SO(2)-equivariant robot policies for tabletop manipulation tasks. In comparison to prior works, our proposed architecture is equivariant to position, orientation, and uniform scaling. In addition, our method can be applied in various 3D manipulation tasks that involve rigid, deformable, and articulated objects.
Appendix B Ablations and Analysis
Ablations on architecture designs. We perform more ablation experiments to understand how various aspects of our proposed method contribute to the final performance of the method. The first question we want to answer is how much the architecture of our method matters compared to that of the prior work EquivAct [16]. To show this, we provide a comparison of EquiBot with a naive extension of EquivAct with a diffusion head. As seen in Figure 8, EquiBot significantly outperforms this architecture for all hyperparameter variations we tried in the OOD setups. When naively adding a diffusion head to EquivAct with action horizon 1, the policy’s poor performance can be explained by its inability to generate multiple actions in a sequence to ensure temporal action consistency [8]. When we increase prediction horizons, the EquivAct architecture, which is not designed to deal with large prediction horizons, suffers from dramatic drops in performance. Our method performs the best across all comparisons we performed in Figure 8.
Figure 9: Ablations on equivariance in object covering. We show that translation, rotation, and scale equivariance each contribute to our method.
Ablations on equivariance. To show how different forms of equivariance play a role in the final performance of our method, we perform ablation experiments in the Object Covering task by subtracting rotation, translation, and scaling equivariance implementations from our method. The result of this ablation experiment is shown in Figure 9. As shown in the results, when we subtract out any form of equivariance from the implementation, the performance of this ablated method suffers from performance drops when the evaluation setup is out-of-distribution to the equivariance this method supports.
Figure 10: Training stability.
Training stability. In the real world, it is impractical to test multiple training checkpoints on the robot. Therefore, high training stability can make checkpoint selection and getting robust test-time performance easy. Here, we specifically compare the training stability of our method with prior work EquivAct [16]. In Figure 10, we plot the in-distribution performance of EquivAct against our method in the Push T task during the training process. We plot average reward over 40 evaluation episodes for every 50 epochs of training. Our method achieves more stable training performance across checkpoints than EquivAct.
Appendix C Further Evaluations and Performance Analysis
C.1 Failure Analysis